OpenStudy (anonymous):

The height of the tunnel at the center is 58 ft and the vertical clearance must be 29 ft at a point 21 ft from the center. Find an equation for the ellipse. Could someone walk me through how to create an equation for this?

OpenStudy (amistre64):



OpenStudy (amistre64):

start with the basic equation of an ellipse with an under y of 58

OpenStudy (amistre64):

\[\frac{x^2}{a^2}+\frac{y^2}{58}=1\] let x=21 and y = 29 to solve for a^2

OpenStudy (amistre64):

might have to be 58^2 under the y :)

OpenStudy (anonymous):

Ahh that makes sense. Thank you! I never understood you could plug them back in.

OpenStudy (amistre64):

\[\frac{x^2}{a^2}+\frac{y^2}{58^2}=1\] \[\frac{21^2}{a^2}+\frac{29^2}{58^2}=1\] solve for a^2 \[\frac{21^2}{a^2}=1-\frac{29^2}{58^2}\] \[\Large \frac{21^2}{1-\frac{29^2}{58^2}}=a^2\]